OpenStudy (anonymous):

How do you find the equation of a plane through the origin and is parallel to another plane with equation 4x-2y+7z+12=0.

OpenStudy (anonymous):

I can use point normal formula? since point is origin so ax+by+cz=0

OpenStudy (anonymous):

But how do I get the a, b and c here?

OpenStudy (anonymous):

That is the normal?

OpenStudy (anonymous):

reuired parallel plane is 4x-2y+7z+k=0 since it passes through (0,0,0) we havve k=0 hence the required plane is 4x-2y+7z=0

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

I studied that planes are parallel if and only if their normal vectors are prallel . For normal vectors to be parallel they need to be the scalar multiples of each other?

OpenStudy (anonymous):

So why 4, -2 and 7? They can be the scalar multiples of each other too?

OpenStudy (anonymous):

@Zeerak plane parallel to ax+by+cz+d=0 is given by ax+by+cz+k=0 where k is to be determined by some other given condition and the one ur saying is valid here

OpenStudy (anonymous):

I dont understand apologies :S

OpenStudy (anonymous):

Can you please say that again?

OpenStudy (anonymous):

@Zeerak as u quoted "I can use point normal formula? since point is origin so ax+by+cz=0 and But how do I get the a, b and c here" so a,b,c is proportional to 4,-2,7 i.e a/4 =b/-2=c/7 ....(as the planes are parallel) i hope now u can ...